Please someone help me with these. I am not sure I am doing them right.
Solve:
1. y square = 49
The square root of 49 is _7 and + 7
y = -7 and + 7
2. (z + 3) squared = 36
z squared + 9 = 36
z squared= 36-9=27= (3x3x3)
=============================
~~~~~~~I can't go on....please help me
3. x squared + 18x - 17 = 0
x^2+18x-17=0
delta= (18)^2-4(1)(17)=256
square root of delta will be:16
therefore we will have two solutions
x= (-19-16)/2 = -35/2
y= (-19+16)/2= -3/2
Thank you very much. Carrie

Found 2 solutions by stanbon, curioustoknow:
Answer by stanbon(75887)

Solve:
1. y square = 49
The square root of 49 is -7 or + 7
y = -7 or + 7
------------------------------------
2. (z + 3) squared = 36
The square root of 36 is -6 or +6
z+3 = 6 or z+3 = -6
z = 3 or z = -9
---------------------------------
3. x^2 + 18x - 17 = 0
This problem requires use of the Quadratic Formula.
If you are familiar with it, use it to get:
x = [-18 +- sqrt(18^2-4*1*-17)]/2
---
x = [-18 +- sqrt(392)]/2
---
x = [-18 +- 19.8]/2
---
x = 1.8/2 or -37.8/2
---
Final Ans: x = 0.9 or x = -18.9
=======================================
Cheers,
Stan H.

Answer by curioustoknow(5)

Please someone help me with these. I am not sure I am doing them right.
Solve:
1. y square = 49
The square root of 49 is _7 and + 7
y = -7 and + 7
2. (z + 3) squared = 36
z squared + 9 = 36
z squared= 36-9=27= (3x3x3)
=============================
~~~~~~~I can't go on....please help me
3. x squared + 18x - 17 = 0
x^2+18x-17=0
delta= (18)^2-4(1)(17)=256
square root of delta will be:16
therefore we will have two solutions
x= (-19-16)/2 = -35/2
y= (-19+16)/2= -3/2
Thank you very much. Carrie